# Describing Globally Distributed Software Architectures for Tax Compliance

### v2

@@ 4.2.4. Identifying jurisdictions of component owners
4.3. Architecture view
In this section, we discuss the practical view construction using the presented viewpoint at a large multinational enterprise.
In detail, we use microservices as the most suitable abstraction layer for software components for the specific software architecture in our case. A microservice represents a discrete piece of a larger application’s functionality and operates as a self-contained unit. Microservices are typically designed to be lightweight, modular, and independently scalable, allowing for greater flexibility and agility in software development (Newman, 2021). Therefore, microservices meet the definition of a software component.
@@ 4.3. Architecture view
In our case company, a component owner is responsible and accountable for its software component. That implies ownership includes functional and non-functional requirements, such as security, reliability, and quality, and extends throughout the software lifecycle, including development, testing, and maintenance. It also involves all aspects of operating that code in production, such as deployment, monitoring, troubleshooting, and remediation of operational and security incidents.
In our case company, each microservice is usually owned by a single team. We excluded 42 microservices (1.64%), each owned by one individual, because the locations of individuals were not accessible for this research due to privacy concerns. Those remaining 2518 microservices have 16533 dependencies between them and are owned by 336 different teams.
@@
Due to the large amount of data, we present the architecture view in two aggregated visualizations instead of a plain list, as graph and as table.
The directed graph as in Figure 3 depicts jurisdictions (countries) as vertices and dependencies as directed edges. An edge goes from the component owner to the component user. We aggregate the number of dependencies between countries as edge weights.
@@
Figure 3. The geographical distribution of the microservice architecture of our case company using our architectural viewpoint. The arrows indicate the direction of the dependency from the microservice user to the microservice owner. The countries are partially substituted by equivalent countries to ensure confidentiality for the company.
@@
Table 2. An equivalent tabular representation of Figure 3 but including unknown dependencies.
@@
5. Evaluation results
In this section, we report the evaluation results from the panel of tax experts (see Section 3 for more details) evaluating the software architecture description with its viewpoint that framed the concerns of tax authorities and the view of software architecture from our industry partner.
5.1. Viewpoint
@@ 5.2. View
5.3. Software architecture description
@@ 6. Discussion
7. Limitations
In this section, we critically reflect on the two components of our simulation-based approach: the software architecture description from our industry partner, the evaluation of the software architecture description in the contrived setting, and the required anonymity.
7.1. Software architecture description
We created the view as described in the paper. We relied on historical information extracted from an internal system that extracts the software product’s component structure (which also comprises microservices) from the code repositories. We believe that this system is a reliable and trustworthy data source for extracting the data we need since the data is also used internally for other purposes. The internal team responsible for maintaining the system helped us to avoid misinterpretation of the data. The date of extraction was randomly selected from the first two quarters in 2023. Although we did not collect this data, we assume that the microservice architecture does not change at a daily rate but is stable, at least for over weeks.
7.2. Contrived setting
The evaluation of the software architecture description for tax compliance highly depends on our tax panel. The contrived setting and the strict anonymization of the case company required us to remove context and potentially relevant information, which includes the specific business model, the revenue stream, and the organizational structure of our case company. Not being able to provide that information limits the generalizability of our evaluation. However, we believe that we provided enough context and information for the panel to provide solid and high-level feedback for the early stages of our research.
@@ 7.2. Contrived setting
The interviews were held in English and German. To avoid translation mistakes and misunderstandings, the second participating interviewer checked the findings after coding and the final version of the paper was sent to all interviewees, asking for confirmation.
7.3. Anonymization
To maintain strict anonymity for the company, we decided not to include participants from the company. The author team spoke on their behalf. We believe that in this early stage of the software architecture description, the author team who extracted the data brought enough context to the table. In future work addressing the subsequent questions of valuing software components, the involvement of internals, however, will become crucial.
We believe that anonymization has a substantial influence on the quality of the evaluation, though. We removed (intentionally) the context of the software architecture. Of course, this context also plays a crucial role in taxation since software engineering does not happen in the void. In particular, the economic context could not be revealed, which led to several loose ends during the evaluation. We believe that this is a major limitation of our findings since the devil is in the details. We could not investigate the potential effects of economic or managerial decisions on the software architecture.
8. Conclusion
We are torn for the conclusion of this research.
@@ 8. Conclusion
On the one hand, our novel software architecture description for reporting globally distributed software architecture for tax compliance reveals that almost 70% of all reuse relationships between the 2560 microservices in our case are cross-border and, therefore, taxable. The extent is magnitudes higher than what we would have expected. Therefore, we urge multinational enterprises to start collecting information about their software architecture and ownership continuously. Accurately collecting this information is never trivial but nearly impossible in hindsight for complex software systems. At this point, we would like to emphasize that the quality and rigor of the data curation at our case company is extraordinary and, from our experience, not industry standard. Not all companies can rely on such a solid foundation grounded in historical and accurate data on their software architectures. This uncertainty introduces further risks and lets tax audits stray away from an empirical-based, data-driven discussion on the value of software. The panel of tax experts found that our proposed architectural viewpoint properly and sufficiently frames the concerns of taxation stakeholders. Although the resulting view of the globally distributed software architecture of our industry partner has to cope with a significant amount of uncertainty, tax experts are used to working with uncertainty (there is no black or white in tax audits but a lot of gray, as interviewee R summed up) and appreciate the evidence-based, empirical approach for visualizing the software architectures to tax experts as software-engineering laymen.
On the other hand, however, we were surprised to be confronted by more fundamental questions that still remain unanswered in software engineering (research). Those questions touch both, currently insufficient definitions of core concepts (such as code ownership or software component), but also limitations in our ways of doing research itself. Although we believe that our research laid the foundation for a data-driven approach to tax compliance of globally distributed software architectures, we believe that tax compliance in software engineering requires us, the community of interested researchers, to take a step backwards first. In the following subsections, we discuss questions we believe are important to address first and the avenue towards this new frontier of taxing software engineering.
8.1. What is code ownership?
The probably most critical and likely most complex issue to solve is our insufficient definition and understanding of code ownership, which needs to be rethought and extended for taxation purposes. As we learned in the course of this study (see Section 5), the definition of ownership in the context of software engineering and the definition of ownership in taxation overlap but are not congruent due to the different contexts and perspectives towards ownership.
@@ 8.1. What is code ownership?
Ownership in software engineering, on the one hand, refers to the assignment of responsibility and accountability for specific portions of software to individuals or teams within a development organization. There are different perspectives on ownership and autonomy in software development organizations. Software development organizations usually adopt the weak ownership principle (Fowler, 2006). This principle entails assigning ownership of specific components (or microservices) to individual teams. In an ideal scenario, the designated team assumes responsibility as the primary (or exclusive) contributor to the associated component. Nevertheless, empirical analysis of real-world instances often reveals a departure from this idealized model (Borg et al., 2023; Tanveer et al., 2023; Greiler et al., 2015; Bird et al., 2011). In practice, a given component may offer support to multiple business streams, undergo development by distinct teams, and necessitate the involvement of specialized teams to address non-functional requirements, such as security and reliability. Consequently, while a specific team retains responsibility for the component, the extent of its contribution to the component can exhibit considerable variability (Zabardast et al., 2022).
Thongtanunam et al. argues that code ownership comes with responsibility, and developers who author the majority of changes to a module are presumably the owners (Thongtanunam et al., 2016). However, contribution may come in other forms, e.g., in modern code review by criticizing code changes authored by other developers, and the study was based on open-source projects. We also believe that results found in open-source projects cannot be directly transferred without further consideration due to the differences in governance structures, incentives, and economic mechanics.
Tax professionals, on the other hand, begin from a civil law perspective and understand ownership in terms of the legal or economic owner regarding the right of exploitation (OECD, 2022). While the identification of the legal or economic owner is relevant for balance sheet allocations, in the context of transfer pricing, the legal owner is considered (only) as the starting point for a functional and risk analysis that aims for a profit allocation commensurate with the value-adding activities performed by the legal entities within the group, the so-called functional ownership (OECD, 2022).
We believe that describing and discussing the intersection, as well as the difference in definitions, requires further more interdisciplinary research. This work is not per-se a purely academic exercise but an endeavour with significant practical relevance for tax compliance.
8.2. What is a software component?
Although the definition of software components are vague and context-specific (for example, (Broy et al., 1998)), the hierarchical, tree-like nature of components and, therefore, software architectures, makes us worrying that we the level of abstraction (i.e. microservices in this research) is the optimal choice.
First, it goes without saying, that also non-microservice software architectures are subject to tax compliance.
Second, our architecture description is, like all models, a simplified abstraction of reality. This simplified abstraction enables us to understand the reality or, in our case, derive taxable conclusions from the software architecture. No model fits all purposes and can, therefore, be perfect, and all are prone to error. This model error is a result of the superposition of two error curves: (i) the bias error, which results from too simple models and can be mitigated by increasing the model complexity, and (ii) the variance error, which reduces the uncertainty with growing model complexity (see Figure 4).
[…]
Figure 4. The overall model error is a result from the superposition of two error curves: (i) the bias error, which results from too simple models and can be mitigated by increasing the model complexity, and (ii) the variance error, which reduces the uncertainty with growing model complexity. We are unsure if we hit the optimal model complexity with our software architecture description.
For our work, this means that we might have abstract software architecture on a too-high-level understanding of software components or a too-low level of software components. On the one hand, high-level understandings such as microservices or even systems (which would then lack a clear definition again) leave more—potentially relevant—information unmodelled but leave the model complexity reasonable. For example, we use only one unspecific relationship between components: the use relationship, which neglects the detailed characteristics of a use relationship or the tree-like nature of software components, resulting in more than 2500 microservices. On the other hand, a low-level understanding of software components (for example, on file level) results in very complex and gigantic models describing the software architecture. In such a software architecture description, the inherent uncertainty would be even more salient.
Therefore, we are unsure if our software architecture description is prone to underfitting or overfitting.
8.3. How can we report empirical evidence?
Many research endeavours that emerge from academia-industry collaborations are prone to the challenge that not all data can be shared “as-is” for confidentiality reasons. We would normally argue for, at least, partial disclosure, anonymization of disclosed data, or sharing of metadata (and, where reasonable, making data conditionally available upon request).
For this research line, however, absolute anonymity and confidentiality between the researchers and all involved parties are paramount to foster trust, honesty, and candor among the participants to ensure that individuals freely contribute highly sensitive information without fear of identification or reprisal. While only an absolute dedication to the anonymity of our case and all involved subjects can safeguard confidentiality and privacy, it also poses a natural barrier to any data dissemination and, more generally, to open science, which we ourselves hold dear as the de-facto standard for empirical software engineering (Mendez et al., 2020).
@@ 8.3. How can we report empirical evidence?
In our line of research, however, we are simply not able to reveal any details about the company, the software architecture, or the interviews without potentially revealing the identity of the involved parties. Even how we collected the data or what data source we used would pose the risk, even if remotely, of revealing the identity of the participating company. This constitutes a fundamental barrier to reproducibility, also in light of that we deem a replication with publicly available data not feasible nor ethical (also considering that we can also not make a case for using open source projects which we discussed already to be not comparable). At the same time, we may well argue that our reporting is still suitable to uncover the potential of describing globally distributed software architectures for the purpose of critically analyzing tax compliance and, at the same time, more fundamental questions that open an interesting avenue towards this new frontier of taxing software engineering. In any case, this general dilemma still remains, in our view, something we deem important to discuss in the research community of empirical software engineering.
@@
Due to the nature of the research, supporting data, except interview guidelines [^2] https://zenodo.org/records/10214942, is not available. Please see our discussion in Section 8.3.
Acknowledgements.

### v3

@@ 4.2.4 Identifying jurisdictions of component owners
4.3 Architecture View
In this section, we discuss the practical view construction using the presented viewpoint at a large multinational enterprise. Understandably, we are not able to describe the case and the data collection in detail to maintain the anonymity of our case company, which was used as a source for creating our view. However, we may well argue that our case company is exemplary for a multinational enterprise developing software. In brief, to create the view, we extracted the software architecture and the team information from an internal system that extracts the component structure of the system from the code repositories on a daily basis.
We use microservices as the most suitable abstraction layer for software components for the specific software architecture in our case. A microservice represents a discrete piece of a larger application’s functionality and operates as a self-contained unit. Microservices are typically designed to be lightweight, modular, and independently scalable, allowing for greater flexibility and agility in software development (Newman, 2021). Therefore, microservices meet the definition of a software component.
@@ 4.3 Architecture View
In our case company, a component owner is responsible and accountable for its software component which includes functional and non-functional requirements, such as security, reliability, and quality, and extends throughout the software lifecycle, including development, testing, and maintenance. It also involves all aspects of operating that code in production, such as deployment, monitoring, troubleshooting, and remediation of operational and security incidents.
In our case company, each microservice is usually owned by a single team. We excluded 42 microservices (1.64%) each owned by a single person, because the locations of individuals were not accessible for this research due to privacy concerns. Those remaining 2518 microservices have 16533 dependencies between them and are owned by 336 different teams.
@@
Due to the large amount of data, we present the architecture view in two aggregated visualizations, as a graph (Figure 3) and as a table Table 2. The directed graph depicts jurisdictions (countries) as vertices and dependencies as directed edges. An edge goes from the component user to the component owner. We aggregate the number of dependencies between countries as edge weights.
@@
Figure 3: The geographical distribution of the microservice architecture of our case company using our architectural viewpoint. The arrows indicate the dependency of the microservice user to the microservice owner. The countries are partially substituted by equivalent countries to ensure confidentiality for the company.
@@
Table 2: An equivalent tabular representation of Figure 3 but including unknown dependencies.
@@
5 Evaluation Results
In this section, we report the evaluation results from the panel of tax experts (see Section 3 for more details) evaluating the software architecture description with its viewpoint that frames the concerns of tax authorities and the view of software architecture from our industry partner.
5.1 Viewpoint
@@ 5.2 View
5.3 Software Architecture Description
@@ 6 Discussion
7 Limitations
In this section, we critically discuss the limitations of our study and their impact on our findings.
7.1 Anonymization
The strict anonymization of the case company required us to remove context and potentially relevant information, which included the specific business model, the revenue stream, and the organizational structure of our case company. Not being able to provide that information limits the generalizability of our evaluation. However, we believe that we provided enough context and information for the panel to provide solid and high-level feedback for the early stages of our research.
To maintain strict anonymity for the company, no participant from the company participated in the panel. The author team spoke on behalf of the company. We believe that in this early stage of the software architecture description, the author team who worked intensively with the company in other research projects and extracted the data brought enough context to the table. In future work addressing the subsequent questions of valuing software components, the involvement of internals, however, will become crucial.
We believe that anonymization has a substantial influence on the quality of the evaluation, though. We removed (intentionally) the economic context of the software architecture. Of course, this context also plays a crucial role in taxation since software engineering does not happen in the void. In particular, the economic context could not be revealed, which led to several loose ends during the evaluation. We believe that this is a major limitation of our findings since the devil is in the details. We could not investigate the potential effects of economic or managerial decisions on the software architecture.
7.2 Data Source for Software Architecture
For creating the software architecture description, we relied on historical information extracted from an internal system that continuously collects the software product’s component structure (which also comprises microservices) from the code repositories. We believe that this system is a reliable and trustworthy data source for extracting the data we need since the data is also used internally for other purposes. The internal team responsible for maintaining the system helped us to avoid misinterpretation of the data. The date of extraction was randomly selected from the first two quarters in 2023. Although we did not collect this data, we assume that the microservice architecture does not change at a daily rate but is stable, at least for over weeks.
7.3 Panel Composition
@@
8 Conclusion
The panel of tax experts found that our proposed architectural viewpoint properly and sufficiently frames the concerns of taxation stakeholders. Although the resulting view of the globally distributed software architecture of our industry partner has to cope with a significant amount of uncertainty, tax experts are used to working with uncertainty (there is no black or white in tax audits but a lot of gray, as interviewee R summed up) and appreciate the evidence-based, empirical approach for visualizing the software architectures to tax experts as software-engineering laymen. However, we believe that we only see the tip of the iceberg of tax compliance challenges for software-developing companies.
@@ 8 Conclusion
First, we found almost 70% of all reuse relationships between the 2560 microservices in our case are cross-border and, therefore, taxable. The extent is magnitudes higher than what we would have expected. Therefore, we urge multinational enterprises to start collecting information about their software architecture and ownership continuously. Accurately collecting this information is never trivial but nearly impossible in hindsight for complex software systems. At this point, we would like to emphasize that the quality and rigor of the data curation at our case company is extraordinary and, from our experience, not industry standard. Not all companies can rely on such a solid foundation grounded in historical and accurate data on their software architectures. This uncertainty introduces further risks and lets tax audits stray away from an empirical-based, data-driven discussion on the value of software.
Second, it goes without saying, that also non-microservice architectures are subject to tax compliance. Future research is needed to discuss the challenges of other architectural styles. This may require a broader discussion that defines a software component that goes beyond the existing vague and context-specific definitions of software components (for example, (Broy et al., 1998)).
@@
Third, as we learned in the course of this study (see Section 5), the definition of ownership in the context of software engineering and the definition of ownership in taxation overlap but are not congruent due to the different contexts and perspectives towards ownership. Ownership in software engineering, on the one hand, refers to the assignment of responsibility and accountability for specific portions of software to individuals or teams within a development organization. There are different perspectives on ownership and autonomy in software development organizations. Software development organizations usually adopt the weak ownership principle (Fowler, 2006). This principle entails assigning ownership of specific components (or microservices) to individual teams. In an ideal scenario, the designated team assumes responsibility as the primary (or exclusive) contributor to the associated component. Nevertheless, empirical analysis of real-world instances often reveals a departure from this idealized model (Borg et al., 2023; Tanveer et al., 2023; Greiler et al., 2015; Bird et al., 2011). In practice, a given component may offer support to multiple business streams, undergo development by distinct teams, and necessitate the involvement of specialized teams to address non-functional requirements, such as security and reliability. Consequently, while a specific team retains responsibility for the component, the extent of its contribution to the component can exhibit considerable variability (Zabardast et al., 2022). Tax professionals, on the other hand, begin from a civil law perspective and understand ownership in terms of the legal or economic owner regarding the right of exploitation (OECD, 2022). While the identification of the legal or economic owner is relevant for balance sheet allocations, in the context of transfer pricing, the legal owner is considered (only) as the starting point for a functional and risk analysis that aims for a profit allocation commensurate with the value-adding activities performed by the legal entities within the group, the so-called functional ownership (OECD, 2022). We believe that describing and discussing the intersection, as well as the difference in definitions, requires further interdisciplinary research. This work is not per se a purely academic exercise but an endeavor with significant practical relevance for tax compliance.
@@
And, last but not least, many research endeavors that emerge from academia-industry collaborations are prone to the challenge that not all data can be shared “as-is” for confidentiality reasons. We would normally argue for, at least, partial disclosure, anonymization of disclosed data, or sharing of metadata (and, where reasonable, making data conditionally available upon request). For this research line, however, absolute anonymity and confidentiality between the researchers and all involved parties are paramount to foster trust, honesty, and candor among the participants to ensure that individuals freely contribute highly sensitive information without fear of identification or reprisal. While only an absolute dedication to the anonymity of our case and all involved subjects can safeguard confidentiality and privacy, it also poses a natural barrier to any data dissemination and, more generally, to open science, which we ourselves hold dear as the de-facto standard for empirical software engineering (Mendez et al., 2020). In our line of research, however, we are simply not able to reveal any details about the company, the software architecture, or the interviews without potentially revealing the identity of the involved parties. Even how we collected the data or what data source we used would pose the risk, even if remotely, of revealing the identity of the participating company. This constitutes a fundamental barrier to reproducibility, also in light of that we deem a replication with publicly available data not feasible nor ethical (also considering that we can also not make a case for using open source projects which we discussed already to be not comparable). At the same time, we may well argue that our reporting is still suitable to uncover the potential of describing globally distributed software architectures for the purpose of critically analyzing tax compliance and, at the same time, more fundamental questions that open an interesting avenue toward tax compliance in software engineering. In any case, this general dilemma still remains, in our view, something we deem important to discuss in the research community of empirical software engineering.
@@
Due to the nature of the research, supporting data, except interview guidelines [^2] https://zenodo.org/records/10214942, is not available. Please see our discussion in Section 8.
Conflict of Interest
The authors declare that they have no conflict of interest.
Acknowledgement